Spreadsheet basics hwk Saved to this PC References Mailings Review View Help Design Layout Tell me what you want to do A Aa- A,ッ. A. 닐를: 仨. a- E L Normal ITNo Spac Heading 1 Heading 2 T Paragraph Styles Z. Determne the numencal results that you would expect to be returned by the formulas in the tollowin spreadsheet. Type your answers in the column to the right of the formulas. 2 Simon Pie Co. Weekly Profit Model 4 Input Data Parameters 5 Selling Price per Pie $10 35.060.0 8 Unit Cost,Fruit Filling (S per pie) 1.00 9 Unit Cost Dough (S per pie) 10 Fixed Cost (Sper week) -HB 13 <35. 1.75.205) 10 12 Decision variables: 14 ·Assure all made can be sold 16 Outputs- Financial results S per 13 Pies made sold per week 30 week enue 18 20 Total ingredients cost 21 Fixed costs 22 Total Cost 24 Total Profit -B17-B22
Answer #1
column C
1
2
3
4
5
6
7 0.40 [because B13<35
8
9 1.75 [because B13<35
10
11
12
13
14
15
16
17 30*10=300
18
19 30*0.40=12
20 30*(1+1.75)=82.5
21 10
22 12+82.5+10=104.5
23
24 300-104.5=195.5

please revert if have any doubts